OpenStudy (anonymous):

Derivative. Is this correct? f(x)=arcsin(x/cos(x)) (formatting and work in reply)

OpenStudy (anonymous):

\[f(x)=\arcsin(\frac{ x }{ \cos(x) })\] \[f'(x)=\frac{ 1 }{ \sqrt{1+(\frac{ x }{ \cos(x) }})^2 }(\frac{ \cos(x)-x(-\sin(x)) }{ \cos(x)^2 })\] I didn't simplify much, so it's easier to see first if I'm doing it right.

OpenStudy (anonymous):

whatever you are doing, i have a suggestion as ugly as this sadistic problem is going to be, it is going to be easier to start with \[\arcsin(x\sec(x))\]

OpenStudy (anonymous):

only mistake i see is that you should have a minus sign inside the radical

OpenStudy (anonymous):

\[f'(x)=\frac{ 1 }{ \sqrt{1-(\frac{ x }{ \cos(x) }})^2 }(\frac{ \cos(x)-x(-\sin(x)) }{ \cos(x)^2 })\]

OpenStudy (anonymous):

it would make it a lot easier to write if you use \(x\sec(x)\) the answer will be the same though you are correct

OpenStudy (anonymous):

ah yeah, that should be minus. ok good. I double checked with wolfram, but it simplifies to the max and it looked so different, I wasn't sure if I was doing it right. http://www.wolframalpha.com/input/?i=derivative+of+arcsin%28x%2Fcos%28x%29%29

OpenStudy (anonymous):

i don't really think it is a matter of taking your answer (which is correct) and using a bunch of trig formulas if you replace \(\frac{x}{\cos(x)}\) by \(x\sec(x)\) you can use the product rule rather than the quotient rule

OpenStudy (anonymous):

the first part is the same, the second would be \[\sec(x)+x\sec(x)\tan(x)\]

OpenStudy (anonymous):

I see, because sec is 1/cos. So x/cos would be x sec (x).

OpenStudy (anonymous):

then as wolfram did you can factor out the \(\sec(x)\) so that the numerator will be \[\sec(x)(\tan(x)+1)\]

OpenStudy (anonymous):

it is just that the product rule comes out a lot neater than the quotient rule, but again everything you wrote (except the plus sign ) is correct

OpenStudy (anonymous):

I never would have thought that it would be easier dealing with sec than cos. I just worked it out with sec and see that yeah, it is a lot neater because of being able to use that product rule instead of the quotient rule. Thanks!